Question 5: A point charge of -100 nC is surrounded by oil having a permittivity of 20 x 10-12 C2N 1m2. How much work would be done on a proton in transporting it from a radial distance of 0.50 m to 1.50 m? Additional challenges: Question 6: Two very small metal spheres carrying charges of 60 uC and -25 uC are located at coordinates (0,0) and (3 m, 0) respectively, in air. How much work would have to be done to bring a third sphere with a charge of -10 uC from very far away to the point (0, 4 m)? Question 7
Answer #1

Question 5)

Potential due to the charge q-100 nC at distance rı 0.50 m is

4TTET1

Potential due to the charge q-100 nC at distance 1.50 m is

V_2=rac{q}{4piepsilon r_2}

V_2-V_1=rac{q}{4piepsilon }left ( rac{1}{r_2}-rac{1}{r_1} ight )

Work done in moving a proton from rı 0.50 m to 1.50 m is W=e(V_2-V_1)=rac{qe}{4piepsilon }left ( rac{1}{r_2}-rac{1}{r_1} ight )

-100 1091.6 1019/1 (20 * 10-12) 1.50 0.50

11.-8.5 * 10-11 J

---------------------------

Question 6)

Charge q_1=60,mu C located at y(0,0) m

Charge q_2=-25,mu C located at (2.2, y2) = (3.0) m ​​​​​​​

Third charge q_3=-10,mu C located at (x_3,y_3)=(0,4),m

Work done on the third charge to bring it from infinity to point (x_3,y_3)=(0,4),m is

4213 193

60*10-6 *-10 * 10-6 -25 * 10-0 *-10 * 10-o 4π(8.85 * 10-12 ) V/(0)2+ (4)2 4π(8.85 * 10-12)V-3)2+ (4)2

W=-1.3485+0.4495=-0.9,J